OpenStudy (anonymous):

what is the distance between the following points (-2sqrt7,10)(4 sqrt 7, 8)

OpenStudy (campbell_st):

its simply a case of substitution into the distance formula \[d = \sqrt{4\sqrt{7} + 2\sqrt{7})^2 + (8 - 10)^2}\] now simplify and calculate...

OpenStudy (anonymous):

d = sqrt(18+4 sqrt(7))

OpenStudy (campbell_st):

so the problem then becomes \[d = \sqrt{(6\sqrt{7})^2 + (-2)^2}\]

OpenStudy (anonymous):

so is it just 256

OpenStudy (campbell_st):

well you need to take the square root of 256

OpenStudy (anonymous):

so 16

OpenStudy (campbell_st):

that's correct